# Hierarchical Clustering Algorithms on Poisson and Cox Point Processes

### v2

@@ 4 Hierarchical Clustroid Nearest Neighbor Chain Algorithm on Cox Point Processes
Specifically, the first example (Example 4.2) is constructed so that each realization contains distinct aggregates of points (see Remark 4.1), and aims at showing how the algorithm forms clusters when such Cox aggregates are present. The second example (Example 4.3) generalizes the first one by removing the requirement for distinct aggregates, while still only retaining a sub-point process of the PPP. As a result, the properties established in Section 2 (such as the finiteness of clusters at each hierarchical level and the weak convergence of the limiting graph) continue to hold. Moreover, it is shown in these examples how the $\mathrm{CHN}^{2}$ clustering algorithm effectively detects the presence of aggregates.
Remark 4.1.
@@
Example 4.2 (Cox point process).
@@
A point process $N$ on a state space $\mathcal{S}$ is called a Cox point process driven by the (locally finite) random measure $M$ if, conditioned on $M$, $N$ is a Poisson point process with intensity function $M$ (see [10]).
Consider a stationary point process $\xi$, and let $r=\{r_{x}\}_{x\in\xi}$ be a real-valued random sequence of marks representing the radii associated with each point in $\xi$. Define ${\Xi}=\bigcup_{x\in\xi}B_{r_{x}}(x)$, where $B_{r_{x}}(x)$ denotes the ball of radius $r_{x}$ centered at $x$. Next, define the random measure $M$ on $\mathbb{R}^{d}$ by
$M(B)=\mu(B\cap\tilde{\xi}),$
where $\mu$ is the Lebesgue measure on $\mathbb{R}^{d}$, and $B\subseteq\mathbb{R}^{d}$ is a Borel set.
Let $N$ be the Cox point process driven by $M$. Conditioned on the realization of $\xi$, $N$ is a sub point process of a homogeneous Poisson point process. More precisely, $N$ is a Cox point process with intensity measure $M$, supported only within the random region $\tilde{\xi}$. Furthermore, due to the stationarity of $\xi$, viewed as a marked point process, $N$ is also stationary.
Since $N$ is a sub-point process of a homogeneous Poisson point process, it does not admit any second-order descending chain. Additionally, the stationarity of $N$ ensures that all results from Section 2 apply to $N$.
Figures 11 and 12 provide simulations of the $\mathrm{CHN}^{2}$ clustering algorithm on $N$, where $N$ is defined as above.
Figure 11: A realization of the Cox point process from Example 4.2, where $S$ contains three random points in $\mathbb{R}^{2}$ and $r=\{40,20,30\}$. This realization consists of approximately 2000 points, which merge into a single cluster after six steps.
Figure 12: A realization of the Cox point process from Example 4.2, where $S$ contains four random points in $\mathbb{R}^{2}$ and $r=\{30,20,30,30\}$. This realization consists of approximately 2500 points, which merge into a single cluster after seven steps.
@@ 4.1 Stopping Criteria for Detecting Data Aggregation in $\mathrm{CHN}^{2}$ Clustering
To answer this, the mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm are computed and plotted for each data set in Figure 13. In this figure, the left diagram corresponds to the data set of Figure 11, while the right one corresponds to the data set of Figure 12. In both diagrams, one can observe a distinct ’jump’ in the mean distances, which can be used to determine the stopping level for clustering.
@@
The results of this method are illustrated in Figure 14, where significant jumps in the mean distance ratios are observed at specific levels. For the dataset in Figure 11, the first pronounced jump occurs at level $6$, while for Figure 12, it appears at level $5$. This is for a threshold $\tau=0.3$.
Figure 13: Left Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, serving as a null hypothesis and the data set shown in Figure 11. At level 6, a pronounced jump is observed in the Cox case, suggesting visually a reasonable stopping criterion. A smaller jump is also visible at level 5, which may indicate partial detection of early-stage clusters.
Right Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, and the data set shown in Figure 12. The first significant jump is observed at level 5, suggesting visually a suitable stopping level for this aggregated dataset.
Figure 14: Left Diagram: Ratio of mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the Cox dataset shown in Figure 11. Clustering is detected at level 6 for a threshold $\tau=0.3$.
Right Diagram: Ratio of mean distances at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the dataset shown in Figure 12. Clustering is detected at level 5 for a threshold $\tau=0.3$.
The following example is a more general case of Example 4.2.
Example 4.3.
Consider $\hat{S}$, an arbitrarily defined random subset of $\mathbb{R}^{d}$. Similar to Example 4.2, define a random measure $M$ on $\mathbb{R}^{d}$ as:
$M(B)=\mu(B\cap\hat{S}),$
where $\mu$ is the Lebesgue measure on $\mathbb{R}^{d}$, and $B\subseteq\mathbb{R}^{d}$ is a Borel set.
Let $N$ be the Cox point process driven by $M$. Conditioned on the realization of $\hat{S}$, $N$ is a homogeneous Poisson point process restricted to the random set $\hat{S}$. So it does not admit any second-order descending chain. Therefore, all the results from Section 2 apply to $N$.
5 Proofs
5.1 Proof of $\mathcal{F/F}$ property of the pre-limits point-shift graphs of $\mathrm{CHN}^{2}$
@@
Remark 5.2.
A descending chain $\{x_{n}\}_{n\in\mathbb{N}}$ in a point process $N$ is defined by the property that for all $i\geq 1$, the inequality $d_{i}<d_{i-1}$ holds, where $d_{i}=|x_{i+1}-x_{i}|$ (see [22] and [11]). Any descending chain in $N$ is a second-order descending chain; however, the converse does not hold.
@@
By classical arguments, it is enough to show that if $\Phi$ is a homogeneous PPP, the probability of having a second-order descending chain in $\Phi+\delta_{0}$, starting from 00, is zero. Consider the following events:
$A=\{\text{ there exists a second-order chain, $\{x_{n}\}_{n\in\mathbb{N}}$, in$\Phi+\delta_{0}$, with $x_{0}=0$}\},$
@@
$A_{R}=\{\text{ there exists a second-order chain, $\{x_{n}\}_{n\in\mathbb{N}}$, in $\Phi+\delta_{0}$, with $x_{0}=0$, and $d_{0},d_{1}<R$}\},$
where $R$ is an integer. Since $A=\cup_{R=1}^{\infty}A_{R}$, it is enough to show that $\mathbb{P}(A_{R})=0$ for all $R$. For $n\geq 3$, let $A_{R,n}$ be the event of existence of a second-order descending chain, $\{x_{i}\}_{i=0}^{n}$ of length (number of edges) $n$, that is, a chain in $\Phi+\delta_{0}$, with $x_{0}=0$, satisfying the property $d_{i}<\max(d_{i-1},d_{i-2})$ for all $12\leq i\leq n-$ 1, and such that $d_{0},d_{1}<R$. Then one can write $A_{R}=\cap_{n\geq 3}A_{R,n}$. Since for all $i$, $A_{R,i}\supseteq A_{R,i+1}$, so
@@
The definition of the second-order descending chain, together with the fact that we consider chains without repeated vertices and the stationarity of $\Phi$, imply that the last equation is equal to
@@
$\mathbb{E}[X_{R,n}]=\frac{\bigl{(}\lambda^{2}\,\omega_{d}^{2}\,R^{2d}\bigr{)}^{\lfloor n/2\rfloor}}{\lfloor n/2\rfloor!}\;\times\;\bigl{(}\lambda\,\omega_{d}\,R^{d}\bigr{)}^{\,n\bmod 2}.$ (7)
@@
For even numbers the base case is $n=0$. The $\mathbb{E}[X_{R,0}]$, is the number of second-order descending chain with length 00 in $\Phi\cup{0}$, with $x_{0}=0$, which is equal to 1, so the base case holds for the even numbers. Assume that (7) holds for an even number $n$. It then follows from (6) that
@@
[…]
@@
where ${\mathbb{S}}^{d}$ denotes the $d$-sphere. Equation (5.1) together with (5.1) give
@@
[…]
@@
[…]
@@
    Figure 16: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on $e_{i}$ is a $d^{0}$-type edge in the infinite path of the $f^{1}$-graph, if it exists.
If $e_{i}$ is a $\delta$-type edge, then there are three possibilities for $e_{i-1}$ and $e_{i-2}$.
@@
    both are $\delta$-type edges (see Figure 17.a). In this case, $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)$ since these edges are in a descending chain in $\{S^{0}_{i}\}_{i}$.
@@
Figure 17: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on $e_{i}$ is a $d$-type edge in the infinite path of the $f^{1}$-graph if it exists.
Propositions 5.3 and 5.6 give

### v3

@@ 3 Clustroid Hierarchical Nearest Neighbor Chain Algorithm on Cox Point Processes
Specifically, the first example (Example 1) is constructed so that each realization contains distinct aggregates of points (see Remark 5), and aims at showing how the algorithm forms clusters when such Cox aggregates are present. As a result, the properties established in Section 2 (such as the finiteness of clusters at each hierarchical level and the weak convergence of the limiting graph,) continue to hold, see Theorem 3.1. Moreover, it is shown in this example how the $\mathrm{CHN}^{2}$ clustering algorithm effectively detects the presence of aggregates. see Subsection 3.1.
Remark 5.
@@
Example 1 (Cox point process).
@@
A point process $N$ on a state space $\mathcal{S}$ is called a Cox point process (driven by the random measure $M$) if, conditioned on $M$, it is a Poisson point process with intensity measure $M$ (see, e.g., [Bartek, Section 2.3]). Equivalently, $N$ is a Poisson point process whose intensity measure is itself random.
Let $\Xi\subseteq\mathbb{R}^{d}$ be a random closed subset. Define the random measure $M$ on $\mathbb{R}^{d}$ by
$M(B)=\mu(B\cap\Xi),\qquad B\in\mathcal{B}(\mathbb{R}^{d}),$
where $\mu$ denotes the Lebesgue measure on $\mathbb{R}^{d}$. Let $N$ be the Cox point process driven by $M$. Conditioned on the realization of $\Xi$, $N$ is a homogeneous Poisson point process restricted to the random region $\Xi$; in particular, it does not admit any second-order descending chain. If the random set $\Xi$ is stationary in law (for example, generated from a stationary marked point process), then $N$ is also stationary.
Special case. In the simulations of Figures 6(a) and 6(b), the random set $\Xi$ is chosen as a union of random balls, $\Xi=\bigcup_{x\in\xi}B_{r_{x}}(x)$, where $\xi$ is a stationary point process and $\{r_{x}\}_{x\in\xi}$ are i.i.d. random radii. This construction illustrates the behavior of the $\mathrm{CHN}^{2}$ clustering algorithm when applied to Cox point processes exhibiting spatial aggregation.
The following theorem shows that all properties established for the homogeneous PPP extend to this Cox setting.
Theorem 3.1.
Let $N$ be a stationary point process in $\mathbb{R}^{d}$ that almost surely admits no second-order descending chain (Def. 4.1), and assume the same general-position (no ties) conditions as in Section 2. This includes, in particular, any stationary sub–point process of a homogeneous PPP and the Cox point process defined in Example 1.
Then the following properties hold for $\mathrm{CHN}^{2}$ on $N$:
  1. (i)
    For each $n\in\mathbb{N}$, the connected components of the $f^{n}$-graph are almost surely in the $\mathcal{F/F}$ class (finite with a unique cycle).
  2. (ii)
    The sequence of $f^{n}$-graphs converges locally weakly to a unimodular limit graph.
  3. (iii)
    Under the Palm version of $N$, the connected component of the root in the limiting graph belongs to the $\mathcal{I/I}$ class (almost surely one-ended with all foils infinite).
Moreover, if there exists a nontrivial bijective point-shift on $N$ (so that $[G,o]$ and $[G,v]$ have the same distribution for every vertex $v$), then property (iii) holds for all components; that is, every connected component of the limit is $\mathcal{I/I}$ almost surely.
Proof.
(i) Finiteness of the pre-limit graphs. The proof follows by repeating the argument of Theorem 2.8 with $\Phi^{0}$ replaced by $N$. That proof only uses (a) stationarity and (b) the absence of (second-order) descending chains to rule out infinite paths (via the reduction to second-order chains as in Proposition 4.5). Both hypotheses hold here by assumption, hence every $f^{n}$-component is a.s. $\mathcal{F/F}$
(ii) Unimodular local weak limit. The proof follows by the proof of Corollary 2.7: the vanishing intensity of cyclic points across levels implies local stabilization of the $f^{n}$-graphs on bounded windows, yielding a local weak limit $[G,0]$; unimodularity of the limit then follows from unimodularity of the prelimits and closedness of unimodularity under local weak limits (see also [Baccelli2017]).
(iii) One-endedness of the component of the root. This follows from adapting the proof of Corollary 2.10 (or Remark 3): under the Palm version of $N$, evaporation holds because the intensity of exit points tends to zero as the level increases. Hence, the descendant tree of the root is finite while the ancestral ray is infinite. By the foil-classification theorem for unimodular networks ([Baccelli2017, Thm. 3.10]), the rooted component is $\mathcal{I/I}$ almost surely.
Upgrading to all components under a bijective point-shift. When a bijective point-shift exists, the distributions of $[G,o]$ and $[G,v]$ are identical for every vertex $v$ (cf. Proposition 2.9 for the PPP case). The $\mathcal{I/I}$ property then extends from the Palm root to every vertex, so all components are $\mathcal{I/I}$ almost surely. In particular, this condition is satisfied by the stationary Cox point process of Example 1, since the existence of a bijective point-shift for any stationary non-equidistant point process follows from Timár’s theorem [timar2009, Theorem 1]. ∎
(a) A realization of the Cox $P$ (Example 1) with $S=\{3$ points}\} and $r=\{40,20,30\}$; $\sim$2000 points; merges after six steps.
(b) A realization of the Cox $P$ (Example 1) with $S=\{4$ points}\} and $r=\{30,20,30,30\}$; $\sim$2500 points; merges after seven steps.
Figure 6: CHN2 on two Cox realizations.
@@ 3.1 Stopping Criteria for Detecting Data Aggregation in $\mathrm{CHN}^{2}$ Clustering
To answer this, the mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm are computed and plotted for each data set in Figure 7. In this figure, the left diagram corresponds to the data set of Figure 6(a), while the right one corresponds to the data set of Figure 6(b). In both diagrams, one can observe a distinct ’jump’ in the mean distances, which can be used to determine the stopping level for clustering.
@@
The results of this method are illustrated in Figure 8, where significant jumps in the mean distance ratios are observed at specific levels. For the dataset in Figure 6(a), the first pronounced jump occurs at level $6$, while for Figure 6(b), it appears at level $5$. This is for a threshold $\tau=0.3$.
Figure 7: Left Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, serving as a null hypothesis and the data set shown in Figure 6(a). At level 6, a pronounced jump is observed in the Cox case, suggesting visually a reasonable stopping criterion. A smaller jump is also visible at level 5, which may indicate partial detection of early-stage clusters.
Right Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, and the data set shown in Figure 6(b). The first significant jump is observed at level 5, suggesting visually a suitable stopping level for this aggregated dataset.
Figure 8: Left Diagram: Ratio of mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the Cox dataset shown in Figure 6(a). Clustering is detected at level 6 for a threshold $\tau=0.3$.
Right Diagram: Ratio of mean distances at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the dataset shown in Figure 6(b). Clustering is detected at level 5 for a threshold $\tau=0.3$.
4 Proof of $\mathcal{F/F}$ property of the pre-limits graphs of $\mathrm{CHN}^{2}$
@@
Remark 6.
A descending chain $\{x_{n}\}_{n\in\mathbb{N}}$ in a point process $N$ is defined by the property that for all $i\geq 1$, the inequality $d_{i}<d_{i-1}$ holds, where $d_{i}=|x_{i+1}-x_{i}|$ (see [Meester] and [Daley]). Any descending chain in $N$ is a second-order descending chain; however, the converse does not hold.
@@
By classical arguments, it is enough to show that if $\Phi$ is a homogeneous PPP, the probability of having a second-order descending chain in $\Phi+\delta_{0}$, starting from 0, is zero. Consider the following events:
[…]
@@
[…]
where $R$ is an integer. Since $A=\cup_{R=1}^{\infty}A_{R}$, it is enough to show that $\mathbb{P}(A_{R})=0$ for all $R$. For $n\geq 3$, let $A_{R,n}$ be the event of existence of a second-order descending chain, $\{x_{i}\}_{i=0}^{n}$ of length (number of edges) $n$, that is, a chain in $\Phi+\delta_{0}$, with $x_{0}=0$, satisfying the property $d_{i}<\max(d_{i-1},d_{i-2})$ for all $2\leq i\leq n-1$, and such that $d_{0},d_{1}<R$. Then one can write $A_{R}=\cap_{n\geq 3}A_{R,n}$. Since for all $i$, $A_{R,i}\supseteq A_{R,i+1}$, so
@@
The definition of second-order descending chains, together with the fact that we consider chains without repeated vertices and the stationarity of $\Phi$, imply that the last equation is equal to
@@
Using this recursive equation, it will be shown in Lemma 4 that the equation (4) holds for $\mathbb{E}[X_{R,n}]$, so when $n$ goes to infinity, the number of descending chain of length $n$ in the Poisson point process goes to zero, and this complete the proof. ∎
@@
$\mathbb{E}[X_{R,n}]=\frac{\bigl(\lambda^{2}\,\omega_{d}^{2}\,R^{2d}\bigr)^{\lfloor n/2\rfloor}}{\lfloor n/2\rfloor!}\;\times\;\bigl(\lambda\,\omega_{d}\,R^{d}\bigr)^{\,n\bmod 2},$ (4)
@@
For even numbers the base case is $n=0$. The $\mathbb{E}[X_{R,0}]$, is the number of second-order descending chain with length 0 in $\Phi\cup{0}$, with $x_{0}=0$, which is equal to 1, so the base case holds for the even numbers. Assume that (4) holds for an even number $n$. It then follows from (3) that
@@
$\int_{B_{R}(x)}(\max(\lVert x\rVert,\lVert y-x\rVert))^{nd}{\mathrm{d}}y$ $=\int_{\theta\in{\mathbb{S}}^{d-1}}\int_{0}^{R}(\max(\lVert x\rVert,\lVert r\rVert))^{nd}r^{d-1}drd\theta$
$=\int_{\theta\in{\mathbb{S}}^{d-1}}\left(\,\int_{0}^{\lVert x\rVert}\lVert x\rVert^{nd}r^{d-1}dr+\int_{\lVert x\rVert}^{R}r^{(n+1)d-1}dr\right)\,d\theta$
$=\int_{\theta\in{\mathbb{S}}^{d-1}}\left(\,\frac{R^{(n+1)d}}{(n+1)d}+\frac{\lVert x\rVert^{(n+1)d}\times n}{(n+1)d}\right)\,d\theta$
@@
where ${\mathbb{S}}^{d}$ denotes the $d$-sphere. Equation (4) together with (4) give
@@
The definition of the point-shift $f^{1}$ implies that the $f^{1}$-graph, as a directed graph, contains two distinct types of edges on $\Phi_{c}^{0}$. The first type is that of edges between the nearest point w.r.t. $d^{0}$. These edges will be referred to as $d$-type edges of level 0 (the green edges in Figure 9). The other type is that of edges that connect an exit point, $x$, of level zero to ${\mathrm{NN1}}(x)$. These edges are referred to as $\delta$-type edges of level one (the red edges in Figure 9).
@@
Figure 9: The structure of an infinite path in $\{S^{0}_{i}\}_{i}$ if it exists. The green arrows show $\mathrm{NN0}$} links ($d$ links of order 0), whereas the red arrows show NN1 links between the exit points of $f^{0}$-cycles and their NN1 points ($\delta$ links of order 1.
@@
    Figure 10: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on the fact that $e_{i}$ is a $d^{0}$-type edge in the infinite path of the $f^{1}$-graph, if it exists.
If $e_{i}$ is a $\delta$-type edge, then there are three possibilities for $e_{i-1}$ and $e_{i-2}$:
@@
    both are $\delta$-type edges (see Figure 11.a). In this case, $\|e_{i}\|<\max(\|e_{i-1}\|,\allowbreak\|e_{i-2}\|)$ since these edges are in a descending chain in $\{S^{0}_{i}\}_{i}$.
@@
Figure 11: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on the fact that $e_{i}$ is a $d$-type edge in the infinite path of the $f^{1}$-graph if it exists.
Propositions 4.2 and 4.5 give